A committee of 3 people is selected at random from 4 men and 8 women. What is the probability that the the committee contains both men and women given that the committee is not composed of all men?
Answer #1

number of ways not composed of all men =N(total number of ways)-N(number of ways with all men)

=(12C3)-(4C3)=220-4 =216

number of ways contains both men and women =N(total number of ways)--N(number of ways with all men)--N(number of ways with all women)

=(12C3)-(4C3)-(8C3)=220-4-56 =160

therefore probability that the the committee contains both men and women given that the committee is not composed of all men =160/216 =20/27